### v3

# Foundational Policy Acquisition via Multitask Learning for Motor Skill Generation

Abstract
In this study, we propose a multitask reinforcement learning algorithm for foundational policy acquisition to generate novel motor skills. Inspired by human sensorimotor adaptation mechanisms, we aim to train encoder-decoder networks that can be commonly used to learn novel motor skills in a single movement category. To train the policy network, we develop the multitask reinforcement learning method, where the policy needs to cope with changes in goals or environments with different reward functions or physical parameters of the environment in dynamic movement generation tasks. Here, as a concrete task, we evaluated the proposed method with the ball heading task using a monopod robot model. The results showed that the proposed method could adapt to novel target positions or inexperienced ball restitution coefficients. Furthermore, we demonstrated that the acquired foundational policy network originally learned for heading motion, can be used to generate an entirely new overhead kicking skill.
Index Terms:
Robot learning; Meta reinforcement learning; Domain Adaptation; Policy selection
publicationid: pubid: $0$–$00$ 2024 IEEE
I Introduction
Multitask reinforcement learning (multitask RL) has been gaining attention as a promising tool to develop human-level skillful agents. In particular, the meta-reinforcement learning (meta-RL) approach that updates shared policy parameters among different policies for different tasks has become a significant research topic for optimizing the parameters of the policies to cope with the multitask RL problems [1, 2, 3, 4]. In this study, we propose to adopt the idea of multitask RL to acquire a foundational motor skill that can adapt to generate novel motor skills by using encoder-decoder networks. As the computational neuroscience studies stated [5, 6, 7], human sensorimotor adaptation mechanisms may contain explicit motor skill learning processes and non-intentional implicit adaptation processes to circumstances. Inspired by these sensorimotor adaptation systems, we propose using a multitask RL framework to acquire a foundational policy that can take the role of implicit adaptation and optimize latent variables of the decoder network to generate novel skills as the explicit motor learning process. Here, we consider the acquisition of a football skill as the target task since football is one of the most popular sports in the world [5]. Furthermore, for example, the ball can fly in completely different directions with a similar heading motion by slightly changing the hitting position, the force applied to the ball, or when the friction of the ball is different. In such cases, continuous task changes may have infinite numbers of task goals, and preparing policy networks for all these tasks becomes impossible. The previous meta-RL framework evaluated the obvious task changes, e.g., the door opening and pick-and-place, and the implicit task change task is insufficient to evaluate and develop.
Therefore, this study proposes a training procedure for policies represented by deep neural networks to cope with the multitask RL problems with the task changes. We were particularly interested in task changes within the same movement category: heading a ball to different goal locations. As the concrete learning framework, we developed a novel skill acquisition method for the contextual Markov decision process (cMDP) to cope with the multitask problems based on a variational inference framework [3]. We demonstrated that selecting a base network model according to the learning performance with randomized task parameters is the key to successful adaptation to multiple reward and environmental settings. Furthermore, our proposed method simultaneously acquires the optimal policy for each context and encoder to embed context variables that correspond to the task settings. The context-related latent variable was estimated to generate skills to handle the novel tasks: unknown rewards and environmental settings. We showed that the Bayesian optimization method can be efficiently used to update the latent variable in the adaptation procedure in cases where the task-relevant variable space is low dimensional (see also Fig. 1).
(a) Multitask heading
(b) Throw in
(c) Hitting ball
Figure 1: Proposed three-phase multitask learning method. The proposed method comprises three phases: acquisition, selection, and generation. In the foundational policy acquisition phase, multiple candidate policies are generated under multiple domains. (a) The $i$-th candidate policy $\pi_{i}$ is trained under the $j$-th domain. In the policy selection phase, one policy is selected from the candidate policy set for the skill generation phase. Low-dimensional latent variables are optimized for adapting to novel tasks and environments in the skill generation phase. (b) The policy selection phase chooses a policy based on the statistical distance from candidate policies $\Pi$. (c) The latent variable $z$ is optimized based on the task performance of the new domain. Multitask heading problem. Monopod robot model tries to acquire policies to hit the ball back to multiple goals in multiple environments. (d, e, f) Implicit multitask heading. Start position, goal position, and the coefficient of the restitution are varied.
We evaluated our proposed policy adaptation method using the monopod motor control task. In this task, as depicted in Fig. 1d-f, the monopod model tries to hit the ball back to the target point while maintaining the posture until it hits the ball. To show the adaptation performance of the proposed method, the target points, as well as the coefficient of restitution that defines the physical interaction between the ball and the monopod model, were changed. We demonstrated that the monopod model was successfully adapted to the implicitly changing task settings. The contributions of our study can be summarized as follows:
  • We developed a policy adaptation method that can cope with implicit multitask problems. The proposed method is composed of three phases: 1) foundational policy acquisition, 2) policy selection, and 3) skill Generation, as shown in Fig. 1a-c.
  • We demonstrated that selecting the base policy from the policy pool derived in the foundational policy acquisition phase in terms of learning performance is the key to successful adaptation.
  • We showed that the monopod model was able to hit the ball back to novel target points with different physical interaction settings using the proposed skill generation method.
The remainder of this paper is organized as follows: In section II, we described the related studies. In section III, we introduced the proposed adaptation method. In section IV, we explained the experimental settings for the monopod model used for evaluating the proposed method. In section V, we presented the results. In section VI, we discussed the comparison of the proposed method with the standard approach as well as the limitations of the proposed method. Finally, in Section VII, we presented the conclusions of the study.
II Related Works
II-A Goal-conditioned RL
To acquire a policy that can cope with multiple task goals, goal-conditioned RL methods have been developed [8, 9, 10, 11, 12]. By acquiring a goal-conditioned value function, the learned policy can generate optimized behaviors according to the continuously changing task goals [9, 13], and [14]. Moreover, some studies proposed to use an image as a task goal [10, 13]. However, these approaches essentially just select a policy corresponding to a specified task goal and cannot handle implicitly changing task goals. This study proposes a three-phase learning and adaptation procedure for coping with implicitly changing environmental settings and task goals.
II-B Data-driven System Identification
To cope with a novel environment by an online data-driven approach, adaptive control methods have been widely studied [15]. If we cannot rely on a parametric environmental model, the model is usually built based on domain-specific knowledge. In that case, we can use an approximated model, possibly represented by a neural network, to bridge the gap between the current estimated and the actual environmental settings. Our early studies proposed collecting data by first using a robust policy that can deal with the novel environment and further improving the policy based on the approximated physical model by using the newly acquired data from the novel environment [16, 17]. Recently, instead of using the robust policy, acquiring a policy network that takes physical parameters as inputs has been proposed [18]. The environment-aware policy was acquired under a variety of environmental settings in the training phase. In the test phase, the physical parameters were identified to improve the performance of the policy further. This approach has been further explored using a low-dimensional representation of the physical parameters. In the training phase, the low-dimensional latent parameter space was extracted. Then, in the test phase, the latent parameters were identified through interaction with the novel environment. The above approach was adopted to deal with the simulation-to-real gap [19, 4]. In this study, we applied these system identification approaches to the implicit multitask RL problem. We showed that the policies can be adapted not only to novel environments but also to novel tasks.
II-C Policy Selection
In this study, we proposed an approach to find an adaptable base policy for unknown tasks and environments. We must consider the policy selection problem [20] to detect the adaptable policy. Our proposed method selected the adaptable base policy by evaluating the learning performances of each policy generated in the foundational policy acquisition phase.
II-D Derivative-free Optimization
For solving optimization problems that have many low-quality local minima, using a gradient-based optimization method is unsuitable. A derivative-free approach such as the evolution strategy can be helpful in such cases. Furthermore, the evolution strategy has been explored for solving optimization problems [21, 22]. In the adaptation phase of our proposed method, we adopted the derivative-free RL approach to avoid low-quality local minima in the context-related variable space extracted by the acquired encoder network. As the derivative-free optimization method, we used the tree-structured Parzen estimator (TPE) [23, 24], which is well known as the high-performance parameter optimization method. In general, derivative-free methods do not work well in high-dimensional optimization problems. However, since the context-related variable space extracted in the adaptation phase was low-dimensional, our proposed method performed well with the derivative-free optimization method.
III Method
In this section, we first defined the cMDP as multitask reinforcement learning, in which the task goal or environments change implicitly. We then introduced the novel variational inference-based reinforcement learning method to cope with the cMDP.
III-A Solving MDP through KL-divergence minimization
Here, we considered a cumulative reward to search for the optimal policy in the maximum entropy RL framework as follows:
$\max_{\pi}\ \mathbb{E}_{\tau|\pi}\left[\sum_{t=0}^{T}\left(r(s_{t},a_{t})-\alpha\log\pi(a_{t}|s_{t})\right)\right],$ (1)
where $\tau=\{s_{0},a_{0},...,s_{t},a_{t},...,s_{T},a_{T}\}$ is the state and action trajectory; $r(s_{t},a_{t})$ is the reward function; $\pi(a_{t}|s_{t})$ is the policy; and the hyperparameter $\alpha\geq 0$ regulates the trade-off between exploration and exploitation[25]. It is known that the entropy-regularized RL problem can be reformulated as the inference problem of the policy distribution, where policy distribution is derived by minimizing the KL divergence between the distributions of the trajectory generated by the policy as,
$q(\tau)=\prod_{t=0}^{T}q(s_{t+1}|s_{t},a_{t})\pi(a_{t}|s_{t}),$ (2)
and the trajectory distribution induced by the reward as,
$p(\tau)=\prod_{t=0}^{T}p(s_{t+1}|s_{t},a_{t})\exp(r(s_{t},a_{t})/\alpha).$ (3)
We then found that minimizing the KL divergence leads to the maximization of the entropy-regularized cumulative reward as follows:
$\alpha\mathrm{KL}[q|p]$ $=\alpha\mathbb{E}_{q}[\log q-\log p]$
$=\mathbb{E}_{q}\left[\sum_{t=0}^{T}\alpha\log\pi_{t}-r_{t}\right].$ (4)
Notably, we can regard the above approach as discounted cumulative reward optimization by simply considering modified dynamics ${\bar{p}}(s_{t+1}|s_{t},a_{t})=\gamma p(s_{t+1}|s_{t},a_{t})$. In this case, transition with $1-\gamma$ into an absorbing state with zero rewards no matter what actions were taken [26].
III-B Solving cMDP
This subsection introduced the proposed variational inference-based RL method for solving the contextual MDP. A cMDP $\{\mathcal{A},\mathcal{S},\mathcal{C},r,p,p_{0}\}$ has the state space $\mathcal{S}$, action space $\mathcal{A}$, context space $\mathcal{C}$, reward function $r(s,a,c)\in\mathbb{R}$, state transition probability $p(s^{\prime}|s,a,c)\in\mathbb{R}$, and initial distribution of the state and context $p_{0}=p(s_{0})p(c)\in\mathbb{R}$. State transition probability and reward functions are conditioned with a vector $c$ in context space $\mathcal{C}$. To acquire the optimal policy in cMDPs, we trained a policy $\pi(a|s,z)$, which depended on states $s$ and the latent variable $z$. Here, the latent variable $z$ is a context-relevant feature vector sampled from an encoder distribution $q(z|c)$. In other words, the encoder converts the current context variables $c$ to the latent variable $z$.
In this study, we proposed a learning procedure to optimize the policy and encoder for a cMDP.
$\min_{\pi,q}$ $\alpha\mathrm{KL}[q|p]=\alpha\mathbb{E}_{q}\left[\log q-\log p\right],$ (5)
$q(\tau,z,c)$ $=p(c)q(z|c)\prod_{t=0}^{T}p(s_{t+1}|s_{t},a_{t},c)\pi(a_{t}|s_{t},z),$
$p(\tau,z,c)$ $=p(c)\rho(z)\prod_{t=0}^{T}p(s_{t+1}|s_{t},a_{t},c)\exp(r(s_{t},a_{t},c)/\alpha).$
The variational distribution $q$ is the probability of trajectory $\tau$ under context $c\sim p(c)$ and latent variable $z\sim q(z|c)$. Likewise, a generative model $p$ is the target distribution close to the variational distribution $q$; $\rho(z)$ is the prior distribution; and $\alpha$ is the scale parameter. The variational distribution and generative model share the context distribution $p(c)$ and transition distribution $p(s_{t+1}|s_{t},a_{t},c)$. Therefore, cumulative rewards with entropy regularization by canceling out shared distribution were obtained as follows:
$\alpha\mathrm{KL}[q|p]$ $=\alpha\mathbb{E}_{q}[\log q-\log p]$
$=\mathbb{E}_{q}\left[\alpha\log\frac{q(z|c)}{\rho(z)}+\sum_{t=0}^{T}\alpha\log\pi_{t}-r_{t}\right]+\mathrm{const}$
$\propto\mathbb{E}_{s_{0},c}\left[\alpha\mathrm{KL}\left[q(z|c)|\rho(z)\right]-V^{\pi}(s_{0},c)\right],$ (6)
$V^{\pi}(s,c):=\mathbb{E}_{\tau,z|s,c}\left[\sum_{t=0}^{T}r_{t}-\alpha\log\pi_{t}\right].$ (7)
In the policy evaluation step, the value function was computed by solving the soft bellman equation as follows:
$Q^{\pi}(s,a,c)$ $:=r(s,a,c)+\mathbb{E}_{p(s^{\prime}|s,a,c)}[V^{\pi}(s^{\prime},c)],$ (8)
$V^{\pi}(s,c)$ $=\mathbb{E}_{q(z|c)\pi(a|s,z)}[Q^{\pi}-\alpha\log\pi].$ (9)
In the policy improvement step, the policy and encoder were updated to minimize the KL divergence as,
$\min_{q,\pi}\mathbb{E}_{s,c}\left[\beta\mathrm{KL}\left[q(z|c)|\rho(z)\right]-V^{\pi}(s,c)\right].$ (10)
By comparing Eq. (6) with Eq. (10), the policy and encoder minimize the KL divergence between the variational distribution and the generative model.
III-C Learning Algorithm
The proposed multitask learning algorithm is composed of three phases: generation, selection, and adaptation phases, as shown in Fig. 1.
III-C1 Foundation Policy Acquisition
In the foundation-policy acquisition phase, the policy was trained under multiple contexts (see Algorithm 1). In the policy evaluation step, we first updated the approximated action value function $Q_{\phi}$ to reduce the Bellman residual as follows:
$\min_{\phi}\mathbb{E}_{\mathcal{D}}\left(Q(s,a,c;\phi)-r(s,a,c)-\gamma\bar{V}(s^{\prime},c)\right)^{2}.$ (11)
We then updated the policy and the encoder using the KL divergence introduced in (10) as follows:
$\max_{\psi,\eta}\mathbb{E}_{(s,c)\sim\mathcal{D}}[\bar{V}(s,c)]-\beta\mathbb{E}_{c\sim\mathcal{D}}\mathrm{KL}\left[q\|\rho\right],$ (12)
where $\mathcal{D}$ denotes the sample set from the replay buffer, and the state value function $V$ can be derived as follows:
$\bar{V}(s,c)=Q(s,a^{\prime},c;\phi)-\alpha\log\pi(a^{\prime}|s,z^{\prime};\psi),$ (13)
$a^{\prime}\sim\pi(a|s,z^{\prime};\psi),\quad z^{\prime}\sim q(z|c;\eta).$
$\phi$, $\psi$, and $\eta$ represent the parameters of the Q function, the policy, and the encoder, respectively. Fig. 2 shows the architecture of each network.
The prior of the latent variable $\rho(z)$ was derived by a standard multivariate normal distribution. The replay buffer stored the tuple $\{s,a,r,s^{\prime},c\}$.
Figure 2: Network Architecture. Our proposed method utilized three neural networks, namely, policy network, encoder network, and $Q$-network which approximate the action-value function. All networks contain two hidden layers, and the numbers on each block represent the layer size. Inputs to the policy network included state $s$ and latent variable $z$. The encoder network considered the context variable $c$ as the input. The $Q$-network recieved state $s$, action $a$, and context $c$ as inputs.
Algorithm 1 Foundation Policy Acquisition
1:  Input: $\phi_{\{1,2\}},\psi,\eta$
2:  $\phi_{\{1,2\}}^{\mathrm{target}}\leftarrow\phi_{\{1,2\}}$
3:  $\mathcal{D}\leftarrow\emptyset$
4:  for each epoch do
5:     sample domain and encoded parameter: $\qquad z\sim q(z|c;\eta),c\sim p(c)$
6:     for collect state-action steps do
7:        $a_{t}\sim\pi(a|s,z;\psi)$
8:        $s_{t+1}\sim p(s_{t+1}|s_{t},a_{t},c)$
9:        $\mathcal{D}\leftarrow\mathcal{D}\cup\{(s_{t},a_{t},r(s_{t},a_{t},c),s_{t+1},c)\}$
10:     end for
11:     for each gradient step do
12:        update policy $\psi$
13:        update encoder $\eta$
14:        update $Q$-function $\phi_{1,2}$
15:        update temperature $\alpha$
16:        update target $Q$-function $\phi_{1,2}^{\mathrm{target}}$
17:     end for
18:  end for
19:  Output $\phi_{\{1,2\}},\psi,\eta$
III-C2 Selection
One of the key ideas of our proposed method was to select a policy model based on the return values in the foundational policy acquisition phase. As a well-known issue, the task performance of a policy highly depends on an exploration procedure determined by a random seed. Thus, in this paper, we proposed the generation of multiple policy models in the offline RL manner to find a high-quality policy that can adapt to a novel task from the foundational policy models. The detailed selection procedure is explained in Algorithm 2. The average reward over the $k$-th episode, $l_{k}$, was stored in the generated $k$-th policy model. We then selected the policy model based on the KL loss in Eq. (10) by substituting the value function with Eq. (7).
Algorithm 2 Selection Phase
1:  Reset candidate policies $\Pi=\phi$
2:  Setup randomized domain with $\mathrm{nun}_{c}$ contexts $\{c\}$
3:  for iteration $k=1,\dots,K$ do
4:     Train $\pi_{k}$ with contexts $\{c\}$triangleright Generation Phase
5:     Rollout $\pi_{k}$ and collect $R(c):=\sum_{t}{\gamma^{t}(r_{t}-\alpha\log\pi_{t})}$
6:     $l_{k}=\frac{1}{\mathrm{num}_{c}}\sum_{c}\left[\beta\log\frac{q(z|c)}{\rho(z)}-R(c)\right]$ in Eq. (10)
7:     $\Pi\leftarrow\Pi\cup\{\pi_{k},l_{k}\}$
8:  end for
9:  Select $\pi_{s}$ based on sorted $\Pi$ by $\{l_{k}\}$
10:  Deploy $\pi_{s}$ to domain with new context $c^{\prime}$
III-C3 Skill Generation
The encoder network learns the latent variable space, where the context variables are embedded in this latent space. In the skill-generation phase, we only optimized the latent variable such that the policy could cope with a novel domain.
To estimate the latent variable $z$ based on the KL loss under the context $c$, we solve the following optimization problem:
[…]
$J(z)$ $=R(z)-\beta\mathrm{KL}[P(*|z)\|\rho(*)],$ (14)
$R(z)=\sum_{t=0}^{T-1}\gamma^{t}(r_{t}-\alpha\log\pi_{t}(z)).$
Skill-generation encoder $P(*|z)$ is represented with multivariate normalized distribution $\mathcal{N}(z,\mathbf{I})$. As we adopted a one-dimensional latent space, black-box optimization methods were utilized. As a concrete optimization method, we chose the TPE, which has been used as a standard tool for searching hyperparameters [23], because it shows empirically better optimization performance than other types of Bayesian optimization methods, such as the GPs and random search methods [23, 24]. To solve the optimization problem, TPE estimates the conditional distribution $p(z|J)$ from sampled data $\mathcal{D}=\{(z,J)\}$. We adopted the off-the-shelf implementation provided in the hyperparameter optimization framework, Optuna [27]. The mean $z$ of encoder distribution is suggested by TPE for each iteration, as shown in Algorithm 3. The algorithm rollouts the policy based on the suggested latent variable $z_{k}$ to collect the return value $J(z_{k})$ and stores it in the replay buffer. The TPE updates the posterior distribution model for the next suggestion. The rollout will be iterated $k_{\mathrm{max}}$ times. Finally, the algorithm returns the best value $z^{*}$ stored in the replay buffer $\mathcal{D}$.
Algorithm 3 Skill Generation Phase
1:  $\pi\leftarrow$ trained policy in Policy Search phase
2:  $\mathcal{D}\leftarrow\emptyset$
3:  TPE model initialization
4:  for iteration $k=1,\dots,k_{\mathrm{max}}$ do
5:     $z_{k}\leftarrow$ TPE suggests
6:     $J(z_{k})\leftarrow$ Rollout policy $\pi(a|s,z_{k})$ in Eq. (14)
7:     Store $(z_{k},J(z_{k}))$ in $D$
8:     TPE updates model $p(z|J)$
9:  end for
10:  Select $z^{*}$ based on the sorted $\mathcal{D}$ by indices $J_{k}$
IV Experiments
We evaluated the proposed learning method using the monopod heading task, as shown in Fig. 1. Simulated experiments were conducted on the sagittal plane using a dynamics simulator MuJoCo [28]. Through this dynamic movement control task, we showed that the monopod robot model can adaptively accomplish multiple tasks in multiple environments using the proposed method.
IV-A Monopod heading task
In the monopod heading task, the robot model tried to hit a ball back to a specified goal point while maintaining its balance until it hit the ball. As small changes in the movement result in significantly different reward values, precise dynamic movement control is necessary. The reward function for the task is defined as follows:
$r=\exp\left(-\lambda_{1}\|p_{\mathrm{ball}}-p_{\mathrm{targ}}(t,c)\|^{2}\right)-\lambda_{2}\|\tau\|^{2},$ (15)
where $p_{\mathrm{ball}}$ denotes the ball position; $p_{\mathrm{targ}}(t,c)$ is the context-dependent desired ball trajectory; and $\tau$ denotes the joint torque vector. The desired ball trajectory was derived using the motion equation from a desired hit position to a target point within a given ball movement duration. The learning trial was terminated if the ball position deviated from the desired ball trajectory by a certain distance. The reward parameters were set to $\lambda_{1}=10.0$ and $\lambda_{2}=5.0\times 10^{-3}$. Furthermore, we terminated the learning trial when the distance between the head position and the desired hitting position exceeded a certain threshold, primarily because the monopod robot fell to the ground.
The monopod robot model had five joints, each controlled by a proportional servo (P) controller. The state variables of the robot are listed in Table I. Notably, the robot did not observe the goal position, although the task achievement was informed through the reward signal. The action variables were the target joint angles for the P controller at each joint.
TABLE I: Observables
Dimensionality Type
Center of mass $2$ float
Head position $2$ float
Foot position $2$ float
Joint angle $5$ float
Joint angular velocity $5$ float
Ball position $2$ float
Ball velocity $2$ float
One-step before action $5$ float
Total dimensionality $25$
In this task, five context variables were considered, which are presented in Table II. The coefficient of restitution defines a physical parameter of the ball and describes the relationship between ball velocities before and after hitting it. Horizontal and vertical goal position variables define different goal locations, and horizontal and vertical throw-in positions represent different ball throw positions toward the monopod robot.
TABLE II: Context variables
Type Settings
Coefficient of restitution float {small, large}
Horizontal goal position float {near, far}
Vertical goal position float {low, high}
Horizontal throw-in position float {near, far}
Vertical throw-in position vertical-axis float {low, hight}
IV-B Network models
The policy network was implemented using the multilayer neural network model. The network model had two hidden layers, shown in Fig. 2. The input variables to the network included the robot state $s$ and the latent variable of the encoder $z$. The encoder network encoded the context variable $c$ to the latent variable $z$ and implemented a neural network model with two hidden layers. In the policy search phase, the encoder and Q-function networks were trained using actual context variables $c$. Subsequently, in the adaptation phase, the encoder network was not used. The latent variable $z$ was estimated through adaptation trials.
IV-C Foundational policy model
In the foundational policy acquisition phase, we trained the networks under the combination of context variables. In other words, we conducted domain randomization with 32 different combinations of context variables as five context variables with two settings were considered, that is, $322^{5}=$ 32. We repeated this learning process 25 times to generate 25 different domain randomized policies with different random seeds.
IV-D Selecting the foundation policy network
Using the policy selection method described in Algorithm $2$, only one policy was selected from the 25 candidate policies.
IV-E Evaluating the skill generation performance
We evaluated how the monopod model can efficiently generated skills to cope with novel tasks and environments using our proposed method. The parameters representing the novel settings were uniformly sampled from intervals defined in Table II. Concretely, we randomly sampled $32$ settings and fixed these values to evaluate the task performances of each method. We confirmed that these $32$ settings differed from those used in the foundational policy acquisition phase described in Section IV-C. We then compared the proposed approach with the policy acquired using the domain randomization method, which is widely used to cope with an unknown environment.
We then further explored the potential of our proposed approach. We set the goal position in the rear region of the robot, while all training in the foundational policy acquisition phase was done with the goal position in the front region. Thus, here we tried to evaluate the extrapolation performance of our skill generation method.
IV-F Evaluating policy selection performance
We compared the proposed policy selection method with random selection in terms of top-one regret, which has been adopted to evaluate policy models [29]. The top-one regret evaluates the error between the best return value and the actual return. In this study, we compared the return value of the selected policy using the proposed method with that of the best policy that acquired the highest return among the candidate policies as follows:
$\frac{1}{N}\sum_{i=1}^{N}|R_{\mathrm{best}}-R_{\mathrm{selected}}|,$ (16)
where $R_{\mathrm{best}}$ denotes the return value of the best policy, and $R_{\mathrm{selected}}$ denotes that of the selected policy using the proposed method. To reliably evaluate the top-one regret, we used a bootstrap method. Concretely, as described in Section IV-C, we generated $25$ candidate policies. Subsequently, we sampled five policies with replacements from the $25$ candidates. Through the sampling procedure described above, we made 1000 sets of the five policies, that is, $N=1000$.
V Results
V-A Skill Generation Performance
Domain randomization is a frequently used approach to cope with multiple tasks and environments. We first evaluated the acquired monopod heading policies using domain randomization as the baseline. We trained the networks under the combination of context variables. The procedure to train the network was the same as that described in Section IV-C. Notably, the proposed method also trained the encoder and policy networks, while the domain randomization method did not.
Fig. 3 and Fig. 3 show the monopod heading task performances at a novel goal position. As shown in the rightmost panel of Fig. 3, the ball deviated from the goal position when a policy acquired by domain randomization was adopted. Using the proposed method, the ball hit by the monopod successfully tracked the desired trajectory and passed around the goal position, as depicted in Fig. 3. Notably, the monopod model also learned to maintain balance when hitting the ball.
Fig. 3 shows the task performances of the policies trained by the domain randomization (green bar) and the proposed adaptation (red bar). Each box plot was depicted based on $32$ task performances, as described in Section IV-E. Both the domain randomization and generation phase of the proposed method trained the networks for $3\times 10^{6}$ time steps, which approximately corresponds to $45,000$ trials. The skill-generation phase only required additional $k_{\mathrm{max}}=100$ trials to perform novel tasks successfully.
(a) Domain randomization
(b) Proposed
(c) Skill generation performance
Figure 3: Monopod heading performances with the novel goal position. The leftmost snapshot shows the timing of hitting the ball, that is, $t=1.0$ s. The green sphere indicates the goal position. The red dot indicates the desired trajectory. (a) Heading performance by domain randomization. As shown in the rightmost panel, at $t=2.0$ s, the ball deviated from the goal position. (b) Heading performance is performed using the proposed method. The hit ball successfully tracked the desired trajectory and passed around the goal position. Notably, the monopod also learned to maintain its balance with the hit ball. (c) Skill generation performance: With Domain Randomization Multitask heading performance. The domain randomization approach could not achieve a satisfactory multitask performance (green bar). Multiple tasks in multiple environments are accomplished using the proposed method (red bar).
V-B Policy selection performance
Here, we evaluate how the policy model selected based on the proposed method performed better than a randomly selected model. We adopted the top-one regret as the performance measure to compare the selection performances. Section IV-F describes a concrete calculation of the top-one regret.
The top-one regret score of the proposed method was $1.4$, whereas that of the random selection was $5.1$. As the smaller regret showed better performance, we confirmed that the proposed method contributed to selecting a potentially well-performing policy after the adaptation phase.
V-C Novel skill generation performance
In this section, we showed how the skill generation phase of our proposed method modifies a policy to cope with a novel task. We only optimized the one-dimensional latent variable $z$ to improve the task performance for the newly provided task. The optimization procedure of the latent variable based on the TPE is explained in Section III-C3.
We depicted how the proposed skill generation method improved the task performance in Fig. 4. As shown in Fig. 4, the hit ball did not track the desired trajectory and deviated from the goal before the adaptation. To evaluate the task performance before the skill generation, we set the latent variable to zero, that is, $z=0.0$, assuming that we did not have prior knowledge of the novel task before the skill generation. After the skill generation, the optimized skill successfully hit the ball towards the new goal position, as shown in Fig. 4.
We also evaluated the overall performances of the policies before and after the skill generation on $32$ different novel tasks, wherein the tasks were sampled randomly, as described in Section IV-E. Fig. 4 shows the task performances of each policy. Before the skill generation, for most novel tasks, task performances were unsatisfactory. Furthermore, performances were varied. After the skill generation, the generated policy for novel tasks consistently achieved successful heading performances.
Furthermore, when we placed the goal position at the back of the robot, the robot model generated a whole new motor skill to ’kick’ the ball backwards to the goal position instead of heading it. Figure 6 illustrated the generated overhead kick movement to hit the ball to the totally new goal position.
The above results indicate that the proposed low-dimensional skill representation and derivative-free optimisation successfully facilitate skill emergence in novel situations.
(a) Foundational
(b) Generated skill
(c) Performances
Figure 4: Multitask Heading performance on novel tasks. The green sphere indicates the goal position. The red dot indicates the desired trajectory. (a) Foundational policy: the hit ball did not track the desired trajectory and deviated from the goal. (b) Generated skill: the adapted policy successfully hit the ball toward the new goal position. (c) Multitask heading performances: With the foundational policy, task performances were unsatisfactory and varied. With the generated skill, generated skills for novel tasks consistently achieved successful heading performances.
(a) Fine-tuning performance
(b) Skill generation perofrmance
Figure 5: (a)Fine-tuning performance. We found that task performances became worse than those before fine-tuning trials. This result indicated that the policy modal could not adapt to new domains with only a small number of additional trials. (b)The proposed method was compared with the baseline method using context variable information. Interestingly, the proposed skill-generation method showed a better heading performance than the method using context information. This result empirically showed that we extracted a meaningful latent space for the given movement category.
Figure 6: Generated overhead kicking skill. We set the goal position (green) in the rear region of the robot, while all training in the foundational policy acquisition phase was done with the goal position in the front region. Thus, here we tried to evaluate the extrapolation performance of our skill generation method. As the result, the robot model generated a whole new motor skill to ’kick’ the ball backwards to the goal position instead of heading it.
VI Discussion
VI-A Skill generation through latent variable optimization
One may consider that the comparison presented in Fig. 3 is not fair because the proposed method conducted additional $100$ trials for adaptation. Therefore, we conducted fine-tuning trials on the novel domain to observe how the policy trained by domain randomization can improve with additional $100$ learning iterations.
Fig. 5 shows the heading performances with the $32$ randomly selected domains. We found that the task performances became worse than those before the fine-tuning trials. This result indicated that the domain randomization policy model could not adapt to new domains with only a small number of additional trials. In other words, extracting the low-dimensional latent space and optimizing the latent variable were significant procedures for generating a skill to cope with novel tasks and environments.
VI-B Using explicit context information
In this study, we assumed that the policy could not observe the context variable in the adaptation phase as we usually lack information regarding novel domains. However, it would be interesting to investigate how task performance can be improved if the context variable was observable. Here, we explicitly encoded the context variable of $32$ novel domains using the trained encoder in the skill-generation phase. The encoder outputs were forwarded to the policy network as latent variables.
Fig. 5 shows the task performances using explicit context information. Interestingly, the proposed adaptation method showed better heading performances than the method using context information. On the one hand, this is because the context variables for novel domains were not used when the encoder network was trained. On the other hand, even the encoder used in our method was not trained with novel context information. However, our proposed method achieved better performance than the method that considers the context variables as inputs. This result empirically showed that we extracted a meaningful latent space for the given movement category.
VII Conclusion
In this study, we proposed a foundational policy acquisition method via the multitask reinforcement learning approach. We particularly focused on dynamic movement tasks in which small movement changes lead to completely different consequences. We consider the multi-heading task to be concrete using the monopod model. The monopod learned to hit a ball back to multiple target positions in this task. Furthermore, the coefficient of restitution of the ball also varied. In the skill-generation phase of our proposed method, context variables such as goal positions and environmental settings were not provided to the monopod model. Therefore, the learning system needed to adapt to implicitly changed contexts. After the skill acquisition process, we showed that the monopod model accomplished $32$ different heading tasks. For the skill acquisition, we selected one foundational policy among $25$ candidate policies based on the return values acquired in the foundational-policy acquisition phase. These policy acquisition and selection strategies to find foundational policy models were important factors for successful skill generation. We further demonstrated that the robot model successfully generated a whole new motor skill to ’kick’ the ball backward to the goal position instead of heading it.
As we focused on a dynamic movement task that only requires a small change to accomplish different targets in different environmental settings, we could expect that the latent variable which represents different contexts can belong to a low-dimensional space. We empirically found that optimizing a one-dimensional latent variable to cope with different situations can effectively achieve higher return values.
As a future study, we aim to explore the possibility of the proposed method on a multi-agent learning problem. We are interested in observing how our proposed method can modify agent policy according to implicit changes in other agents’ strategies.
Acknowledgments
This study used several opensource software projects, such as PyTorch [30], rlkit [31], dm-control [32], and Optuna [27]. We are grateful to the project contributors for providing valuable machine learning tools. This work was supported by JSPS KAKENHI (Grant number: 19J22987); by project JPNP20006, commissioned by NEDO; by JST Moonshot R&D program (Grant number: JPMJMS223B-3); and by Tateisi Science and Technology Foundation.
Appendix
Learning algorithm
We adopted the target networks, clipped double-$Q$ learning, and entropy autotuning in the policy generation phase. The target network and clipped double-$Q$ learning are the methods for reducing the estimation error of $Q$ functions. For this purpose, four neural networks were trained to approximate a single $Q$ function. Entropy autotuning changes the temperature parameter $\alpha$ that controls the exploration and exploitation trade-off.
Clipped double-$Q$ learning
$Q$-function networks were trained to reduce Bellman residuals, as shown in Eq. (17) as follows:
$\min_{\phi_{i}}\mathbb{E}_{\mathcal{D}}\left(Q(s,a,c;\phi_{i})-r(s,a,c)-\gamma\bar{V}(s^{\prime},c)\right)^{2},$ (17)
$\bar{V}(s,c):=\min_{i\in{1,2}}Q(s,a^{\prime},c;\phi_{i}^{\mathrm{target}})-\alpha\log\pi(a^{\prime}|s,z;\psi),$
$a^{\prime}\sim\pi(a^{\prime}|s,z;\psi),z\sim q(z|c;\eta),\{s,a,s^{\prime},c\}\sim\mathcal{D},$
where $\phi_{i}$ and $\phi_{\mathrm{target}}$ are the $Q$-network parameters; $r$, $\gamma$, and $\alpha$ denote the reward function, discount factor, and temperature parameter, respectively; $\pi$ represents the policy; $q$ represents the encoder; and $\mathcal{D}$ is the state-action pair data stored in a replay buffer. The value function $\bar{V}$ was computed from $Q(s^{\prime},a^{\prime},c)$, where $s^{\prime}$, $a^{\prime}$ and $c$ are the state, action, and context at the next step, respectively. Simply using the above method tends to overestimate the action values; therefor, clipped double-$Q$ Learning was proposed to overcome this overestimation problem [33]. To use this method, we trained two networks and selected the smaller output of the two approximated $Q$ functions to derive the value $\bar{V}$. In practice, $Q$ values at the next step are derived from the target network $\phi_{\mathrm{target}}$ explained below.
Target networks
The target networks were used to stabilize the learning process of the $Q$ function. As we explained in the previous section, the target value $V$ was derived from the $Q$ network, which caused the target value estimation to be unstable [34]. Therefore, as suggested in [35], we updated the target network through moving average as follows:
$\phi_{i}^{\mathrm{target}}\leftarrow\tau\phi_{i}+(1-\tau)\phi_{i}^{\mathrm{target}}.$ (18)
The Polyak’s coefficient $\tau$, which ranges from 0 to 1, was used to adjust the update rate.
Entropy autotuning
To autotune the temperature parameter $\alpha$ of the entropy regularization term, we considered the following optimization problem [36]: $H$ is the conditional entropy of the policy distribution given the state $s$, and $\bar{H}$ is the target entropy set by an experimenter. The target entropy was set to $-1$ times the action dimension $|A|$ as follow [36]:
$\max_{\pi}$ $\mathbb{E}_{\tau\sim\mathcal{D}}\left[\sum_{t}\gamma^{t}r_{t}\right],$ (19)
$\mathrm{s.t.}$ $\mathbb{E}_{s\sim\mathcal{D}}[H(\pi)]<\bar{H}.$ (20)
Although this equation does not explicitly contain the temperature parameter $\alpha$, it can be attributed to the following min-max problem as Lagrange duality when the temperature parameter $\alpha$ is regarded as a Lagrange multiplier:
$\min_{\alpha>0}\max_{\pi}\mathbb{E}_{\tau\sim\mathcal{D}}$ $\left[\sum_{t}\gamma^{t}r_{t}-\alpha(\log\pi_{t}+\bar{H})\right].$ (21)
The min-max problem was solved as:
$\min_{\alpha>0}\mathbb{E}_{(a,s,z)\sim\mathcal{D}}$ $\left[\alpha(-\log\pi(a|s,z)-\bar{H})\right].$ (22)
A heuristic transformation $\alpha\leftarrow\exp(\alpha)$ was performed on the temperature parameter $\alpha$ to guarantee the positive value as practical implementation  [^1] https://github.com/rail-berkeley/softlearning/blob/master/softlearning/algorithms/sac.py#L140. Accessed: July-10, 2023..
Policy gradient and reparametrization trick
We used the reparametrization trick to calculate the gradient of the objective function with respect to policy parameters. The objective function was expressed as follows:
$\max_{\psi,\eta}\mathbb{E}_{(s,c)\sim\mathcal{D}}[\bar{V}(s,c)]-\beta\mathbb{E}_{c\sim\mathcal{D}}\mathrm{KL}\left[q\|\rho\right],$ (23)
$a^{\prime}\sim\pi(a^{\prime}|s,z;\psi),z\sim q(z|c;\eta),\rho(z)=\mathcal{N}(z|\mathbf{0},\sigma\mathbf{I}).$
Here, $\phi$ and $\eta$ are the network parameters of the policy $\pi$ and encoder $q$ networks, respectively; and $\rho(z)$ is the prior distribution of the latent variable.
To derive the gradient, expectation over the policy distribution $\pi$ and encoder distribution $q$ must be evaluated as follows:
$\max_{\psi,\eta}$ $\mathbb{E}_{(s,c)\sim\mathcal{D},q\pi}[Q(s,a,c)-\alpha\log\pi(a|s,z;\psi)]$
$-\beta\mathbb{E}_{c\sim\mathcal{D},q}\left[\log\frac{q(z|c;\eta)}{\rho(z)}\right].$ (24)
By defining the related variables, as shown below, the parameters of the neural network do not depend directly on the distributions.
$a$ $=\tanh(f(s,z;\psi)+\sigma(s,z;\psi)\epsilon),\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I}),$ (25)
$z$ $=g(c;\eta)+\delta\omega,\omega\sim\mathcal{N}(\mathbf{0},\mathbf{I}),$ (26)
$P(a;\psi)=P(\epsilon)\left|\frac{da}{d\epsilon}\right|^{-1}_{\psi},P(z;\eta)=P(\omega)\left|\frac{dz}{d\omega}\right|^{-1}_{\eta}.$ (27)
Here, $\epsilon$ and $\omega$ are the random variables sampled from the standard normal distribution; and $f$, $\sigma$, and $g$ are the functions represented by the neural networks. Now, the objective function can be defined as,
$J(\psi,\eta)=$ $\mathbb{E}_{(s,c)\sim\mathcal{D},\omega,\epsilon}[Q(s,a,c)-\alpha\log\pi(a|s,z;\psi)]$
$-\beta\mathbb{E}_{c\sim\mathcal{D},\omega}\left[\log\frac{q(z|c;\eta)}{\rho(z)}\right],$
$a=a(s,\epsilon,z;\psi),z=z(c,\omega;\eta),$
where the integrated variables, $\epsilon$ and $\omega$, are independent of the network parameters. Therefore, the interchange of the gradient and integral for the policy and encoder parameters, $\psi$ and $\eta$, was allowed:
$\nabla J$ $=\nabla\mathbb{E}_{(s,c)\sim\mathcal{D},\omega,\epsilon}\left[Q-\alpha\log\pi-\beta\log\frac{q}{\rho}\right],$
$=\mathbb{E}_{(s,c)\sim\mathcal{D},\omega,\epsilon}\left[\nabla Q-\alpha\nabla\log\pi-\beta\nabla\log\frac{q}{\rho}\right].$ (28)
Hyperparameters
The parameters used in this study are listed in Table III. Notably, most of the parameters were selected based on those reported in the study by Haarnoja et al.[36].
TABLE III: List of hyperparameters
hyperparameter policy search adaptation
action time period 25ms 25 ms
neural network optimizer Adam Adam
Policy and encoder learning rate $3\times 10^{-4}$
Q learning rate 0.001 -
$\beta$ $e^{-5}$ $e^{-5}$
replay buffer size $10^{6}$ 100
collect step size per training 1000 100
collect step size per evaluation 1000 100
update times per training 1000 10
Polyak’s coefficient $\tau$ 0.01 -
batch size 256 100
discount $\gamma$ 0.99 1.00
optimization algorithm Adam Adam
encoder variance $\delta$ 0.01 0.01
prior variance $\sigma$ 1 1
target entropy $-|\mathcal{A}|$
temperature $\alpha$ 0.01
$k_{\mathrm{max}}$ 100